# Charge Regulation Effect on Nanoparticles Interaction Mediated by Polyelectrolyte

### v1

@@ IV Results and Discussion
Figure 7: Charge regulation (CR) effects on NPs–PE interactions at low higher concentrations ($pIp=pIm=3$) with keeping all other parameters same as in Fig. 3. (a) Top: Simulation snapshot of CR simulation at intersurface separation ($D$) between colloids, $7\ell_{b}$. The color scheme is the same as in Fig. 1. Bottom: Simulation snapshot of a constant charge (CC) simulation at the same $D$. (b) Average charge on NPs and PE as a function of $D$. (c) Total force ($F_{X}$), between NPs for CR and CC as a function of $D$. The insert shows $F_{X}$ at low salt concentration. (d) Adsorbed monomer fraction ($f_{ad}$), as a function of $D$. (e) The radius of gyration ($R_{g}$) of the PE versus $D$. Data averaged over 5 (CR) and 50 (CC) equilibrated configurations.
@@ V Conclusion
The highly versatile nature of charged polymers or polyelectrolytes (PEs) provides enormous opportunities for controlled manipulation of nanoparticle’s (NP) physicochemical properties. The adsorption of PE on the surfaces of NPs plays a critical role in modulating their stability and aggregation behavior in solution. This is crucial for various applications. In this study, we have systematically explored the influence of charge regulation (CR) on the interactions between two NPs coated with base groups in the presence of an oppositely charged single weak polyanion in a monovalent solution. By implementing a hybrid CR Monte Carlo/molecular dynamics (CR-MC/MD) simulation framework, we provide a mechanistic understanding of how dynamically regulated surface charges, in contrast to static constant charge (CC) approximations, impact PE adsorption behavior and modulate interparticle forces.
@@
At low salt concentrations ($pIp=pIm=7$), our results show that CR significantly enhances PE adsorption on NP surfaces. This is also confirmed by the $Rg$ behavior. Importantly, the time evolution of the adsorbed monomer fraction is much faster and higher(2-fold) than that observed for CC. This adsorption strongly suppresses electrostatic repulsion by effectively neutralizing the NP surface charges, resulting in weak net interparticle interactions across a wide range of NPs separations. In the addition of salt, long-range, however weak, attractive interactions between like-charged NPs are obtained. The CC simulations demonstrate clear long-range attractions driven by polymer bridging, with the PE chains adopting extended conformations connecting both NPs. However, the CR scenario promotes localized adsorption, where the PE preferentially collapses onto a single NP, thereby eliminating the entropic and elastic contribution associated with bridging and reducing the net attractive interaction.
@@
The dependence of intermolecular forces on PE chain length was found to be markedly different between these two models. While longer PE chains in the CC simulations enhance bridging and, consequently, attractive interactions, the CR simulations exhibit largely chain-length-independent behavior due to the saturation of available adsorption sites. This highlights the dominant role of surface ionization equilibrium in dictating the interaction landscape in the CR paradigm. We further investigated the role of acid-base dissociation constants by tuning the net ionization parameter, $\Delta pK=pK_{a}+pK_{b}-2pH$. This parameter captures the joint contribution of the dissociable groups on both the PE and the NP surfaces. The results demonstrate that the difference between CR and CC is most significant at intermediate $\Delta pK$ values, where partial ionization allows CR to dynamically respond to local electrostatic environments, leading to stronger adsorption and reduced interparticle forces. Notably, CR consistently exhibits faster and more complete adsorption dynamics, emphasizing its responsiveness and kinetic favorability over the CC approximation. At higher salt concentrations ($pIp=pIm=3$), electrostatic screening reduces the effective interactions in both CR and CC simulations. Under these conditions, the differences in the extent of adsorption and radius of gyration between the two models diminish. However, even in the high-salt regime, charge regulation continues to exhibit greater PE adsorption, although the difference in inter-particle forces becomes less pronounced. This underlines the ion-sensitive nature of charge regulation, where the balance between surface ionization and ionic screening governs both adsorption and colloidal/NPs stability. Our findings suggest that surface-tunable charge regulation can be a strategic tool for stabilizing nanoparticle suspensions under physiological ionic conditions. Future work could explore active tuning of CR by external stimuli such as electric fields or pH gradients.
@@ VI Acknowledgment
P.K.J. acknowledges the financial support from the Science and Engineering Research Board (SERB), Department of Science and Technology (DST), India (Grant Nos. CRG/2022/006365 and ITS/2024/001866) and IIT Jodhpur for a Seed Grant (No. I/SEED/PKJ/20220016). SK acknowledges the financial support from the Science and Engineering Research Board (SERB), Department of Science and Technology (DST), India (grant No. EEQ/2023/000676) and IIT Jodhpur for a research initiation grant (grant no. I/RIG/SNT/20240068).

### v2

@@ IV Results and Discussion
Figure 7: Charge regulation (CR) effects on NPs–PE interactions at low higher concentrations ($pIp=pIm=3$) with keeping all other parameters same as in Fig. 3. (a) Top: Simulation snapshot of CR simulation at intersurface separation ($D$) between colloids, $7\ell_{b}$. The color scheme is the same as in Fig. 1. Bottom: Simulation snapshot of a constant charge (CC) simulation at the same $D$. (b) The average charge on NPs and PE as a function of $D$. (c) The total force ($F_{X}$) between NPs for CR and CC as a function of $D$. The insert shows $F_{X}$ at low salt concentration. (d) Adsorbed monomer fraction ($f_{ad}$), as a function of $D$. (e) The radius of gyration ($R_{g}$) of the PE versus $D$. The averages are calculated over 5 (CR) and 50 ($C$) equilibrated configurations.
Finally, we investigate the electrostatic interaction between PE and NPs at a higher salt concentration ($pIp=pIm=3$), while keeping all other parameters identical to those used in Fig. 3. Simulation snapshots at $D=7\ell_{b}$ are depicted in Fig. 7a, revealing similar structural features to those observed under low salt conditions (see Fig. 3a). However, as illustrated in Fig. 7b, the charge on the NPs is nearly twice the magnitude and opposite in sign to that of the PE, due to the increased availability of free ions participating in surface ionization reactions.
The impact of salt concentration on the interaggregate force is depicted in Fig. 7c. In the CC simulations, polymer bridging persists, but at a shorter range of separation than in the low-salt scenario, which is consistent with previous findings Podgornik et al. (1995). In contrast, the CR simulations show no qualitative change in the force profile with increasing salt concentration, except for a noticeable downward shift in $F_{X}$. This reduction in net force arises from the increased population of counterions (primarily anions) required to neutralize the higher positive surface charge on the NPs. Overall, these observations suggest that, under CC conditions, NP-NP interactions are strongly PE-mediated at low salt concentrations but become progressively screened as the ionic strength increases. In contrast, under CR conditions, the interaggregate force remains dominated by counterions across salt concentrations, reflecting the dynamic charge regulation and strong ion-mediated screening intrinsic to these systems. Furthermore, we observe a significant increase in PE adsorption for both CR and CC simulations at high salt concentrations, as shown in Fig. 7d. Importantly, the difference in adsorption between CR and CC systems is significantly reduced compared to that seen under low-salt conditions (see Fig. 3d), which can be attributed to the higher surface charge on the NPs under these conditions. A comparable trend has also been observed experimentally. Huang et al. Huang et al. (2018) showed that increasing salt concentration shortened the range of interaction forces and reduced distinct charge-regulation effects, although adsorption remained slightly higher than in the constant charge, in agreement with our findings. The adsorption difference is mainly due to the enhanced surface affinity in the CR model, which persists even under strong ionic screening. A similar convergence is evident in the radius of gyration (see Fig. 7e), where the disparity between CR and CC simulations diminishes with increasing salt concentration, indicating more compact PE conformations in both cases.
These results highlight that although the CR and CC systems exhibit clear differences at intermediate separations, especially under low ionic strength, these differences diminish at larger separations and higher salt concentrations. In CR simulations, increased salt primarily leads to a suppression of interaggregate forces due to enhanced screening by counterions. In contrast, CC simulations remain sensitive to salt, with polymer-mediated bridging interactions being progressively diminished. Nevertheless, in both scenarios, higher salt promotes greater PE adsorption and a reduction in $R_{g}$, reflecting more compact and surface-associated polymer configurations.
@@ V Conclusion
The highly versatile nature of charged polymers or polyelectrolytes (PEs) provides enormous opportunities for controlled manipulation of nanoparticles (NPs) physicochemical properties. The adsorption of PE on the surfaces of NPs plays a critical role in modulating their stability and aggregation behavior in electrolyte solution. This is crucial for various applications. In this study, we have systematically explored the influence of charge regulation (CR) on the interactions between two NPs coated with base groups in the presence of an oppositely charged single weak PE in a monovalent solution. We employed a hybrid charge-regulation Monte Carlo/molecular dynamics (CR-MC/MD) framework to investigate PE adsorption and its impact on interparticle forces. In addition, constant-charge (CC) molecular dynamics simulations were carried out to assess the impact of charge regulation.
@@
We find that at low salt concentrations ($pIp=pIm=7$), CR significantly enhances PE adsorption onto NP surfaces, leading to the collapse of the PE chain on one of the NPs. This adsorption strongly suppresses electrostatic repulsion by effectively neutralizing the NP surface charges, resulting in weak net interparticle interactions across a wide range of NPs separations. In contrast, the CC simulation shows long-range attractions driven by polymer bridging, with the PE chains adopting extended conformations connecting both NPs.
Additionally, the dependence of intermolecular forces on PE chain length was found to be markedly different between these two models. In CC simulations, longer PE chains enhance bridging and hence attractive interactions. In CR simulations, the behavior is largely chain-length independent due to saturation of adsorption sites.
We further investigated the role of acid-base dissociation constants by tuning the net ionization parameter, $\Delta pK=pK_{a}+pK_{b}-2pH$. This parameter captures the joint contribution of the dissociable groups on both the PE and the NP surfaces. The results demonstrate that the difference between CR and CC is most significant at intermediate $\Delta pK$ values, where partial ionization allows CR to dynamically respond to local electrostatic environments, leading to stronger adsorption and reduced interparticle forces. Notably, CR consistently exhibits faster and more complete adsorption dynamics, emphasizing its responsiveness and kinetic favorability over the CC approximation.
@@
Finally, we show that at high salt concentration ($pIp=pIm=3$), electrostatic screening reduces the effective interactions in both CR and CC simulations. Under these conditions, the differences in the extent of adsorption and radius of gyration between the two models diminish. However, even in the high-salt regime, charge regulation continues to exhibit greater PE adsorption, although the difference in the inter-particle forces becomes less pronounced. This underlines the ion-sensitive nature of charge regulation, where the balance between surface ionization and ionic screening governs both adsorption and colloidal/NPs stability. Our findings, in line with experimental observations, provide strong validation for our CR-MC/MD framework. These results further indicate that surface-tunable charge regulation can be used as a strategic tool for stabilizing nanoparticle suspensions under physiological ionic conditions. Future work could explore active tuning of CR by external stimuli such as electric fields or pH gradients.
@@ VI Acknowledgment
P.K.J. acknowledges the financial support from the Anusandhan National Research Foundation, India (Grant Nos. CRG/2022/006365 and ITS/2024/001866) and IIT Jodhpur for a Seed Grant (No. I/SEED/PKJ/20220016). SK acknowledges the financial support from the Anusandhan National Research Foundation, India (grant No. EEQ/2023/000676) and IIT Jodhpur for a research initiation grant (grant no. I/RIG/SNT/20240068).